# Unexpected properties of Markov chains with an infinitely countable set of states

### v4

@@ Abstract.
Discrete-time Markov chains; recurrence and transience; birth-and-death processes; continuous-time Markov processes, stochastic calculus
@@ 1. Introduction
Let $\mathscr{P}$ be an irreducible Markov chain with the infinite set of states $\{0,1,\ldots\}$. Let $p_{i,j}^{(n)}$ denote $n$-step transition probabilities from state $i$ to state $j$, $p_{i,j}=p_{i,j}^{(1)}$. Recall that a Markov chain is recurrent, if its initial state 00 is visited infinitely many times with probability 1. It is well-known that the Markov chain $\mathscr{P}$ is recurrent, if and only if $\sum_{n=1}^{\infty}p_{0,0}^{(n)}=\infty$ (see [8]). The application of this last condition is very limited, since the derivation of $p_{0,0}^{(n)}$ is generally combinatorially hard. The known applications are chiefly related to simple random walks and can be found in many textbooks on the theory of Markov processes.
In this paper, we suggest a new result that enables us to verify the condition of recurrence or transience for a relatively wide class of irreducible Markov chains. It turns out that the result presented here is an extension of the well-known one from the theory of birth-and-death processes, but to our best knowledge it is unknown. The proof of the main result is based on the reduction to a continuous time Markov process, and, as a result, we arrive at the equations that describe a birth-and-death process. Then the well-known results of the birth-and-death theory are applied. The suggested approach makes the proof of the main result simpler compared to possible routine analysis that would use direct analytic derivations.
@@
An example of Markov chain, the states of which form connected domain is a Markov chain, where $p_{i,j}>0$ for all $i,j\geq 1$ with $i\neq j$. The examples of Markov chains the states of which do not form connected domain are considered in Section 3.
@@
Assume that the states of the Markov chain $\mathscr{P}$ form connected domain, and $\sum_{j=1}^{\infty}jp_{0,j}<\infty$. Denote $e_{i}^{-}=\sum_{j=1}^{\infty}jp_{i+j-1,i-1}$ and $e_{i}^{+}=\sum_{j=1}^{i+1}jp_{i-j+1,i+1}$, $i\geq 1$, assuming that $e_{i}^{-}<\infty$.Then, the Markov chain $\mathscr{P}$ is recurrent if and only if
@@
The assumptions $\sum_{j=1}^{\infty}jp_{0,j}<\infty$ and $e_{i}^{-}<\infty$, $i\geq 1$, are technical, since the infinite series $\sum_{j=1}^{\infty}jp_{0,j}$ and $\sum_{j=1}^{\infty}jp_{i+j-1,i-1}$ appears explicitly in the proof of the theorem. It is possible to avoid these assumptions using additional arguments. Specifically, one can consider the series of Markov chains, in which the required parameters are assumed to increase to infinity and arrive at the conclusion in the limit case.
The rest of this paper is organized as follows. In Section 2, we prove the theorem. In Section 3, we provide a nontrivial example of recurrent Markov chain demonstrating how the statement of the theorem works. In Section 4, we provide counterexamples to demonstrate that the key assumption of the theorem, that the states of a Markov chain forms connected domain, is important. In Section 5, we discuss our findings.
@@ 2.1. Birth-and-death process
(1) $\left(\begin{matrix}0&1&0&0&0&0&\ldots\\ q_{1}&0&p_{1}&0&0&0&\ldots\\ 0&q_{2}&0&p_{2}&0&0&\ldots\\ 0&0&q_{3}&0&p_{3}&0&\ldots\\ \vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right),\quad q_{i}+p_{i}=1,\quad i\geq 1.$
It describes transition probabilities for a birth-and-death process. It is well-known that the Markov chain $\mathscr{X}$ is recurrent if and only if $\sum_{n=1}^{\infty}\prod_{i=1}^{n}q_{i}/p_{i}=\infty$ (see [6], [4, p. 286]).
@@ 2.2. Simplification of the problem
We start from the following notice related to natural extension of (1). Let $c_{1}$, $c_{2}$,…be a sequence, $0<c_{i}\leq 1$, $i\geq 1$. Consider the Markov chain $\mathscr{Y}$ with the transition probabilities matrix
(2) $\left(\begin{matrix}0&1&0&0&0&0&\ldots\\ c_{1}q_{1}&1-c_{1}&c_{1}p_{1}&0&0&0&\ldots\\ 0&c_{2}q_{2}&1-c_{2}&c_{2}p_{2}&0&0&\ldots\\ 0&0&c_{3}q_{3}&1-c_{3}&c_{3}p_{3}&0&\ldots\\ \vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right).$
@@
2.3. Chapman-Kolmogorov equations
We show that the structure of the equations for the final probabilities can be reduced to the same structure as that of birth-and-death processes. For this purpose, it is convenient to switch from the original Markov chain to the continuous time Markov process as follows. We assume that the times between transitions all are independent and exponentially distributed with mean 1.
First, we give three examples to make our further arguments clearer. The first two examples describe elementary models. The third one describes a natural model and is closer to the formulated result. The further example of Section 3 is based on the model considered in the third example of this section.

### v5

# A new criterion for recurrence of Markov chains with an infinitely countable set of states

@@ Abstract.
Discrete-time Markov chains, recurrence and transience, birth-and-death processes, continuous-time Markov processes, stochastic calculus
@@ 1. Introduction
1.1. Formulation of the main result
Let $\mathscr{P}$ be an irreducible Markov chain with the infinite set of states $\{0,1,\ldots\}$. Let $p_{i,j}^{(n)}$ denote $n$-step transition probabilities from state $i$ to state $j$, $p_{i,j}=p_{i,j}^{(1)}$. Recall that a Markov chain is recurrent, if its initial state 00 is visited infinitely many times with probability 1. It is well-known that the Markov chain $\mathscr{P}$ is recurrent, if and only if $\sum_{n=1}^{\infty}p_{0,0}^{(n)}=\infty$ (see [4, 10]). The application of this last condition is very limited, since the derivation of $p_{0,0}^{(n)}$ is generally combinatorially hard. The known applications are chiefly related to simple random walks and can be found in many textbooks on the theory of Markov processes. The more extended study for the recurrence of Markov chains can be found in [4, pp. 65, 66], where the aforementioned elementary result has been extended. Specifically, it was derived the system of equations that clarified the solution of this problem under quite general conditions. However, the derivation of explicit conditions for recurrence or transience from that system of equation is hard and can be provided for only the particular cases such as the Markov chains of the birth-and-death type, for which the aforementioned system of equations has been further applied in [4, pp. 67–71].
In this paper, we suggest a new result that enables us to verify the condition of recurrence or transience for a relatively wide class of irreducible Markov chains. It turns out that the result presented here is an extension of the well-known one from the theory of birth-and-death processes. Furthermore, for a wide class of the Markov chains, the conditions for recurrence or transience are given in the same terms as those for the birth-and-death chains. In other words, we prove that a Markov chain of the considered class is recurrent if and only if a specifically defined birth-and-death chain is recurrent.
@@ 1.1. Formulation of the main result
A trivial example of a Markov chain, the states of which form connected domain is the birth-and-death type chain, where the diagonal elements of the transition matrix are not necessarily zeros. Another example of a Markov chain, the states of which form connected domain is a Markov chain, where $p_{i,j}>0$ for all $i,j\geq 1$ with $i\neq j$. The examples of Markov chains the states of which do not form connected domain are discussed in Section 3.
@@
Assume that the states of the Markov chain $\mathscr{P}$ form connected domain, and $\sum_{j=1}^{\infty}jp_{0,j}<\infty$. Denote $e_{i}^{-}=\sum_{j=1}^{\infty}jp_{i+j-1,i-1}$ and $e_{i}^{+}=\sum_{j=1}^{i+1}jp_{i-j+1,i+1}$, $i\geq 1$, assuming that $e_{i}^{-}<\infty$. Assume also that $1-p_{i,i}>\epsilon$ for all $i\geq 0$, where $\epsilon$ is some positive value. Then, the Markov chain $\mathscr{P}$ is recurrent if and only if
@@
The assumptions $\sum_{j=1}^{\infty}jp_{0,j}<\infty$ and $e_{i}^{-}<\infty$, $i\geq 1$, are technical, since the infinite series $\sum_{j=1}^{\infty}jp_{0,j}$ and $\sum_{j=1}^{\infty}jp_{i+j-1,i-1}$ appear explicitly in the proof of the theorem. It is possible to avoid these assumptions using additional arguments. Specifically, one can consider the series of Markov chains, in which the required parameters are assumed to increase to infinity and arrive at the conclusion in the limit case.
The well-known corollary from Theorem 1.2 is the condition of recurrence or transience for birth-and-death chains (e.g., [4, pp. 66–71]).
1.2. Methodology of the study
The direct methods for the analysis of Markov chains, such as suggested in [4, pp, 65, 66] for instance, seem to be insufficient for the solution of the problem considered in the present paper. The system of equations obtained in [4, p. 66] can be served as a starting point of the study that may require tedious analytical derivations and finally lead to the equations similar to these being used in the paper.
We decided to follow another way. The analysis of this paper is based on the methods of stochastic calculus and the theory of martingales in continuous time and involves the theory of point processes. For this purpose, the original Markov chains is extended to the Markov processes in continuous time. Such approach has an advantage, since the results for the birth-and-death processes and point processes are widely presented in the literature and widely known, while the similar study based on the discrete time processes seems do not yield any simplifications.
For those continuous time Markov processes, we derive the system of stochastic differential equations. We obtain the equations that describe limit behavior of the process and prove that the conditions for recurrence or transience are the same as those for the Markov processes of the birth-and-death type.
1.3. Outline of the paper
The rest of this paper is organized as follows. In Section 2, we prove the theorem. Our methodology is first demonstrated on elementary Example 2.1 and then developed for a general type Example 2.3 that closely describes the proof in the general case. In Section 3, we provide simple but nontrivial examples of recurrent and transient Markov chains that demonstrate how the statement of the theorem works. In Section 4, we provide counterexamples that demonstrate that the key assumption of the theorem stating that a Markov chain forms a connected domain, is important. In Section 5, we discuss our findings.
@@ 2.1. Birth-and-death process
(2.1) $\left(\begin{matrix}0&1&0&0&0&0&\ldots\\ q_{1}&0&p_{1}&0&0&0&\ldots\\ 0&q_{2}&0&p_{2}&0&0&\ldots\\ 0&0&q_{3}&0&p_{3}&0&\ldots\\ \vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right),\quad q_{i}+p_{i}=1,\quad i\geq 1.$
It describes transition probabilities for a birth-and-death process. It is well-known that the Markov chain $\mathscr{X}$ is recurrent if and only if $\sum_{n=1}^{\infty}\prod_{i=1}^{n}q_{i}/p_{i}=\infty$ (see [7], [4, p. 67–71], [5, p. 286]).
@@ 2.2. Simplification of the problem
We start from the following notice related to the natural extension of (2.1). Let $c_{1}$, $c_{2}$,…be a sequence, $0<\epsilon<c_{i}\leq 1$, $i\geq 1$. Consider the Markov chain $\mathscr{Y}$ with the transition probabilities matrix
(2.2) $\left(\begin{matrix}0&1&0&0&0&0&\ldots\\ c_{1}q_{1}&1-c_{1}&c_{1}p_{1}&0&0&0&\ldots\\ 0&c_{2}q_{2}&1-c_{2}&c_{2}p_{2}&0&0&\ldots\\ 0&0&c_{3}q_{3}&1-c_{3}&c_{3}p_{3}&0&\ldots\\ \vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right).$
@@
2.3. Stochastic Chapman-Kolmogorov equations
The variant of the Chapman-Kolmogorov equations used in this section is a stochastic version of these equations that was originally used in [8] for a specified queueing network of a closed type.
We show that the structure of the equations for the steady state probabilities can be reduced to the same structure as that of the birth-and-death processes. For this purpose, it is convenient to switch from the original Markov chain to the continuous time Markov process as follows. We assume that the times between transitions all are independent and exponentially distributed with mean $1/2$.
First, we give three examples to make our further arguments clearer. The first two examples describe elementary models. The third one describes a natural model and is closer to the formulated result. The further elementary examples of Section 3 are based on our general result.
@@ 2.3. Stochastic Chapman-Kolmogorov equations
Note that the models $\mathscr{A}$ and $\mathscr{B}$ are described in the terms of Markov chain transition probability as follows. The model $\mathscr{A}$ is described by matrix (2.1), where $p_{i}=\lambda_{i}/(\lambda_{i}+\mu_{i})$, $q_{i}=\mu_{i}/(\lambda_{i}+\mu_{i})$. Then for the model $\mathscr{B}$ the matrix of transition probability is
(2.3) $\left(\begin{matrix}0&1&0&0&0&0&0&\ldots\\ c_{1}q_{1}&0&0&\frac{c_{1}p_{1}}{2}&0&0&0&\ldots\\ 0&c_{2}q_{2}&0&0&\frac{c_{2}p_{2}}{2}&0&0&\ldots\\ 0&0&c_{3}q_{3}&0&0&\frac{c_{3}p_{3}}{2}&0&\ldots\\ \vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right),$
with the same values $p_{i}$ and $q_{i}$ and with the normalization constants $c_{i}$ such that the sums in each row are equal to $1$. Note that $c_{i}q_{i}=2\mu_{i}/(\lambda_{i}+2\mu_{i})$, and $c_{i}p_{i}/2=\lambda_{i}/(\lambda_{i}+2\mu_{i})$, $i\geq 1$.